OpenStudy (anonymous):

How do I simplify 2√6 / 2√6 + 1? The answer is 24 - 2√6 / 23?.. How? D:

OpenStudy (asnaseer):

use the identity:\[a^2-b^2=(a+b)(a-b)\]to remove the radical expression in the denominator. so, in this case, multiply numerator and denominator by \(2\sqrt{6}-1\)

OpenStudy (anonymous):

that's what I did but should I be multiply 2 x 2 cand √6 x √6?

OpenStudy (asnaseer):

\[\frac{2\sqrt{6}}{2\sqrt{6}+1}=\frac{2\sqrt{6}}{2\sqrt{6}+1}*\frac{2\sqrt{6}-1}{2\sqrt{6}-1}=\frac{2\sqrt{6}(2\sqrt{6}-1)}{(2\sqrt{6}+1)(2\sqrt{6}-1)}\]\[=\frac{(2*\sqrt{6}*2*\sqrt{6}-2\sqrt{6})}{(2\sqrt{6})^2-1^2}=\frac{4*6-2\sqrt{6}}{4*6-1}=\frac{24-2\sqrt{6}}{23}\]

OpenStudy (anonymous):

Although, I don't understand why the exercise is called "simplify". There is nothing simple about the alternate form.

OpenStudy (asnaseer):

having no radicals in the denominator is considered "simpler"

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

why did you do 1^6?

OpenStudy (anonymous):

I meant 1^2

OpenStudy (asnaseer):

\[a(b+c)=ab+ac\]so:\[2\sqrt{6}(2\sqrt{6}-1)=(2\sqrt{6})*(2\sqrt{6})-(2\sqrt{6})*1\]\[=2*2*\sqrt{6}*\sqrt{6}-2\sqrt{6}=4*6-2\sqrt{6}=24-2\sqrt{6}\]

OpenStudy (anonymous):

you multiplied it by 1, why?

OpenStudy (asnaseer):